SUBMILLIMETER-WAVE SPECTROSCOPY OF SHORT-LIVED SPECIES USING A BWO

\underline{ISAMU MORINO^{a}} and KOICHI M. T. YAMADA, National Institute for Advanced Interdisciplinary Research, Higashi 1-1-4, Tsukuba 305-8562, Japan.

A submillimeter-wave spectrometer using a BWO (Backward Wave Oscillator) has been constructed and applied to measurements for shrot-lived species. The BWO frequency was stabilized by a phase-lock loop, using a Schottkytype harmonic mixer which mixes the radiation from the BWO with that of a millimeter-wave synthesizer^{1}. A part of submillimeter power was reflected by a wire-grid beam-splitter and directed to the absorption cell. The transmitted power was detected by a magnetically-tuned, liquid-helium-cooled InSb bolometer. To test performances of the spectrometer, the rotational spectra of NH_{2}OH, NH_{2}, and CF have been measured in 400 – 500 and 600 – 800 GHz region.

The NH_{2} radical is known to be an important intermediate in astrochemistry and various chemical reaction process. The submillimeter-wave absorption spectrum was observed toward the dust continuum source Sagittarius B2^{2}. Recently, Tonooka et al. have measured the microwave spectrum and determined precise molecular constants^{3}.

NH_{2}OH has been interested because of its large amplitude internal motions. The submillimter-wave rotational spectra in the ground and vibrationally excited OH-torsion state have been observed for the first time.

The accuracy of submillimeter-wave line positions of the CF radical reported previously using a far-infrared laser sideband spectrometer is not enough to perform astronomical searches for this radical. In the present study the submillimeterwave spectra of CF have been measured with same precision as microwave spectroscopy.

aDomestic Research Fellow of Japan Science and Technology Corporation. 1G. Winnewisser, Vib. Spectrosc., 8, 241 (1995); I. Morino et al., J. Mol. Spectrosc., 185, 142 (1997). ^{2}E. F. van Dishoeck et al., Astrophys. J., 416, L83 (1993). ^{3}M. Tonooka et al., J. Chem. Phys., 106, 2563 (1997).